OpenStudy (anonymous):

sqrt of 20/125 ?!?!?!

Directrix (directrix):

2/5

OpenStudy (anonymous):

i got .4 same thing :-)

Directrix (directrix):

20/125 = 4/25 √ (4/25) = 2/5

OpenStudy (anonymous):

thanks!